OpenStudy (anonymous):

Evaluate \[125\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

@thivitaa heyy;)

OpenStudy (thivitaa):

3(125) +2

OpenStudy (anonymous):

which is what?

OpenStudy (anonymous):

i know but that is not the answer in the multiple choice. the question is at the top scroll

OpenStudy (anonymous):

@bahrom7893 help please? :( no one can figure it out

OpenStudy (anonymous):

What's there to evaluate?

OpenStudy (anonymous):

these are the multiple choice A. 15,625 B. 125 C. 20 D. 25

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

I'm confused what you want us to do

OpenStudy (anonymous):

scroll up the question is there.. need to evaluate that number

OpenStudy (anonymous):

There's nothing to evaluate

OpenStudy (anonymous):

thats what the question is, idk

OpenStudy (austinl):

To change it to proper form you do as was stated above. \[125\frac{2}{3} = \frac{(125\times 3) + 2}{3}\]

OpenStudy (anonymous):

so its 125 right?

OpenStudy (austinl):

No.

OpenStudy (austinl):

What is 125*3

OpenStudy (anonymous):

375 but thats not one of my multiple choice answers

OpenStudy (anonymous):

i did 375 divided by 3 and got 125 and that was one of the choices

OpenStudy (austinl):

THAT IS PART OF THE ANSWER.... now add 2 to that.